OpenStudy (rissaunnie):

what is the solution of the system x+y=0 y=2x+6

OpenStudy (anonymous):

Okay, do you know how to do substitution?

OpenStudy (rissaunnie):

no

OpenStudy (anonymous):

well, if you take x+y=0, and express y in terms of, x + y =0 y = -x So then for both equations to have the same solution set, both equations must share the same coordinates. Therefore, for the second equation y = 2x + 6, plug in y = -x into y = 2x + 6 and solve for x. Then once you found x, plug it into any of the two equations and solve for y.